OpenStudy (anonymous):

sec theta = -5 (Give answers to the nearest tenth of a degree)

OpenStudy (lalaly):

\[\sec \theta=\frac{1}{\cos \theta}\]

OpenStudy (lalaly):

so\[\cos \theta= -\frac{1}{5}\]

OpenStudy (lalaly):

\[\cos^{-1}(-\frac{1}{5})=\theta\]

OpenStudy (anonymous):

Thank you very much! I was plugging in \[1/\cos^{-1}(-5) \]which didn't fit the domain.

OpenStudy (lalaly):

:D:D